Two 2.00 mC point charges are located on the x-axis. One is at x=1.00 (a) Determine the electric field on the y-axis at y=0.500 m (b) Calculate the electric force on a -3.00C charge placed on the y-axis at y=0.500 m
1
Expert's answer
2020-11-10T07:09:55-0500

(a) Let's draw a free-body diagram:


Let the first 2.00 mC point charge is located at a distance "x=1.0\ m" from the origin and the second 2.00 mC point charge - at a distance "x=-1.0\ m". Let Q be the point where we need to determine the electric field due to two point charges and let "y=0.5\ m" be the distance to that point. As we know, the electric field directed away from the positive electric charges. So, we can denote both fields due to our positive charges as "E_+". Let's draw the "x"- and "y"-components of electric field "E_+". As we can see from the FBD, "E_{+x}" components will canceled. So, we have only two "E_{+y}" components of the electric field.

We can find the magnitude of the electric field from the formula:


"E_+=k\dfrac{q}{r^2},"

here, "k=8.99\cdot 10^9\ N\dfrac{m^2}{C^2}" is the Coulomb's constant, "q=2\cdot 10^{-6} C" is the charge, "r" is the distance between the charge and point Q.

We can find the distance between the charge and point Q from the Pythagorean theorem:


"r=\sqrt{x^2+y^2}=\sqrt{(1.0\ m)^2+(0.5\ m)^2}=1.12\ m."

Since we deal only with "E_{+y}" components we can multiply E_+ by "sin\theta". We can find angle "\theta" from the geometry:


"sin\theta=\dfrac{y}{r}=\dfrac{0.5\ m}{1.12\ m}."

Finally, we can find the net electric field on the y-axis at y=0.500 m due to two point charges:


"E_{net}=2E_{+y}=2E_+sin\theta=2k\dfrac{q}{r^2}\dfrac{y}{r},""E_{net}=2\cdot 8.99\cdot 10^9\ N\dfrac{m^2}{C^2}\dfrac{2\cdot 10^{-6} C}{(1.12\ m)^2}\dfrac{0.5\ m}{1.12\ m}=1.28\cdot10^4\ \dfrac{N}{C}."

(b) We can calculate the magnitude of the electric force from the formula:


"F=E_{net}q=1.28\cdot10^4\ \dfrac{N}{C}\cdot 3.0\cdot 10^{-6}\ C=3.84\cdot 10^{-2}\ N."

Answer:

(a) "E_{net}=1.28\cdot10^4\ \dfrac{N}{C}."

(b) "F=3.84\cdot 10^{-2}\ N."